OpenStudy (anonymous):

. Solve for y, then find the value of y when given x = –2. 6x = 7 – 4y (1 point) A.-12 B. 7/4 C.19/4 D.24

OpenStudy (anonymous):

HELP!!!!!!!

OpenStudy (anonymous):

6*-2 = 7-4y -12 = 7-4y -19 = -4y y = 19/4